OpenStudy (anonymous):

The Phoenix Suns won 19 more than twice as many games as they lost. The Suns played 82 games. How many wins and losses did the team have?

OpenStudy (stormfire1):

I'm getting 31.5 :P

OpenStudy (stormfire1):

losses

OpenStudy (anonymous):

I got loss = 21 win = 61

OpenStudy (anonymous):

how did you get that answer? What equation did you use in the first place?

OpenStudy (stormfire1):

Maybe I'm reading the question wrong but for me, the equation was 82 = 19 + 2x

OpenStudy (anonymous):

game played = 82 won= 19+2y ; where y is the loss game played = win + loss 82= 19+2y+y 82= 19+3y 3y= 82-19 = 63 y= 63/3 = 21 sub y in equation, then won = 19+ 2*21= 61

OpenStudy (stormfire1):

"19 more than twice as many games as they lost" says to me, 19 + 2*losses

OpenStudy (anonymous):

@cmullins clear for you??

OpenStudy (stormfire1):

I'm sure your answer is right though...I just think it's worded funny.

OpenStudy (anonymous):

yes, little confused

OpenStudy (anonymous):

i got this i figured out what i did wrong. Thank you so much